The auto repair shop of Quality Motor Company uses standards to control the labor time and labor cost in the shop. The standard labor cost for a motor tune-up is given below:

Standard Hours Standard Rate Standard Cost
Motor tune-up 2.50 $30.00 $75.00

The record showing the time spent in the shop last week on motor tune-ups has been misplaced. However, the shop supervisor recalls that 50 tune-ups were completed during the week, and the controller recalls the following variance data relating to tune-ups:

Labor rate variance $ 250 F
Labor spending variance $ 200 U

Required:

1. Determine the number of actual labor-hours spent on tune-ups during the week.

2. Determine the actual hourly rate of pay for tune-ups last week. (Round your answer to 2 decimal places.)
Answer #1

Labor efficiency variance = 200U - 250F = 450 U

(SH - AH) *SR = - 450

(2.50*50 - AH) * 30 = - 450

ACTUAL HOURS = 140 HOURS

​​​​​​​​​​​Labor Rate Variance = (SR-AR) *AH

(30-AR)*140 = 250

ACTUAL RATE = 28.21

COMMENT IF YOU FACE ANY ISSUES